OpenStudy (anonymous):

Find the solutions to 0 ≤ x3 – 6x2 + 8x – 2 by graphing f(x) ≤ x3 – 6x2 + 8x – 2 using technology

OpenStudy (anonymous):

i think x3 is x^3

OpenStudy (paxpolaris):

do you mean graph f(x)=x^3-6x^2+8x-2 paste $f\left(x\right)\le x^3-6x^2+8x-2$ into https://www.desmos.com/calculator to graph

OpenStudy (anonymous):

Thanks for the help but I've figured out what I was doing wrong.